OpenStudy (help!!!!):

4. What is 35% of a number if 12 is 15% of a number? A. 5 B. 12 C. 28 D. 33 E. 62 plz show the equation

OpenStudy (anonymous):

(12/x)=(15/100) will tell you the number, given x find y, in (y/x)=(35/100)

OpenStudy (help!!!!):

why is it 12/x?

OpenStudy (anonymous):

in other words, 12*100/15=x

OpenStudy (help!!!!):

nvm i got it